# Cosmological perturbation theory of primordial compact sources

### v1

@@ 1 Introduction
For all sources that obey this condition, the effect of the cosmological background on gravitational wave propagation simply amounts to a redshift of the GW frequency, the modulation of the GW amplitude by the scale factor and a redshift of the masses of the source [Maggiore:2007ulw, Deruelle:1984hq].
@@
The main formalism currently used for the computation of perturbations of FRLW spacetimes is based on the mathematically elegant SVT decomposition, which allows to fully decouple scalar, vector and tensor modes and write down gauge invariant cosmological perturbations [Bardeen:1980kt, Kodama:1984ziu, Durrer:1993tti]. When combined with inflation and quantum mechanics, cosmological perturbation theory allows to infer the large scale structure of the current universe [Mukhanov:1990me].
Due to the spatial non-locality of the SVT decomposition, perturbations are typically expanded in Fourier modes. Fourier domain is convenient for the study of the fundamental stochastic GW background as the spectrum of perturbations generated during inflation is nearly scale invariant [Planck:2015sxf]. However, when considering individual localized sources, the SVT formalism requires to solve Poisson equations in order to reconstruct the metric perturbation, which is not straightforward. It is convenient instead to derive the perturbation equations in position space as most of the tools derived in post-Newtonian methods in the non-cosmological context are precisely derived in position space using a basis of symmetric trace-free (STF) tensors [Damour:1990gj]. In order to fill this gap, we derive in this paper a formalism that solves for the propagation of GW in FRLW universe without using neither the geometric optics approximation nor the SVT decomposition. Instead, the equations will be decoupled in a basis of symmetric trace-free (STF) tensors by introducing a generalized harmonic gauge. Our formalism is therefore adapted to the study of compact sources in the early universe, such as e.g. cosmic strings or primordial black hole mergers [Sasaki:2018dmp]. The generalized harmonic gauge that we define is the first gauge to our knowledge that allows to fully decouple the linearized perturbations without using a SVT decomposition, which departs from standard treatments of cosmological perturbation theory [Baumann:2009ds].
@@
One important question in cosmology is the very definition of a compact source. We emphasize in our treatment that defining a vanishing stress-energy tensor perturbation outside a compact support is fundamentally inconsistent with the local conservation of the stress-energy tensor. This will lead us to define a nearly spatially compact source, which still admits a non-compact energy density perturbation. We will finally derive the explicit metric perturbations for a nearly compact source in power law cosmologies. We will follow the consistent quadrupolar truncation scheme introduced in [Compere:2023ktn].
@@ 2 Decoupled perturbations in generalized harmonic gauge
Let us first specialize to the de Sitter background where $\Lambda=3H^{2}$ and $a=-\frac{1}{H\eta}$, which implies $(\frac{\dot{a}}{a})=-\frac{1}{\eta}$ and q=1q=-1. In that case, the gauge condition reduces to $\bar{\nabla}_{\alpha}\tilde{h}^{\alpha}_{\;\;\mu}=\frac{2\Lambda}{3}\eta\tilde{h}_{0\mu}$, which is identical to the gauge condition defined in [deVega:1998ia]. Around de Sitter, Eq. (2.11) becomes
@@
This equation matches exactly with Eq. (69) of [Date:2015kma], and Eq. (3.23) of [deVega:1998ia] after correctly identifying the field variables. It was shown in [deVega:1998ia] that these linearized equations can be totally decoupled after using an appropriate choice of field variables.
@@
We emphasize that unlike standard cosmological perturbation theory, we have not performed the scalar, vector, tensor (SVT) decomposition of the field variables, which is spatially non-local on constant time sections. Here, we performed a local decomposition into irreducible representations of $SO(3)$, a STF (symmetric tracefree) decomposition. The SVT decomposition in cosmological perturbation theory has found vast applications to non-localized sources in stochastic waves of astrophysical or cosmological origin. Our formalism is suited to solve for inhomogeneous solutions in terms of Green’s functions, as we will develop in Section 3. In particular, one can relate the linearized field in terms of the multipolar structure of spatially compact sources, as performed around Minkowski spacetime [Damour:1990gj].
@@
Importantly, one cannot absorb the right-hand side of Eq. (2.26a) by performing a local shift of the stress-energy tensor. Therefore, in the presence of a varying cosmological fluid variable $a^{2}(p+\epsilon)$, the perturbed stress-energy tensor is not conserved on its own. Using our definition of the shifted stress-energy tensor, this non-conservation can be confined to the $\eta$ component of the shifted stress-energy conservation equation, and be can limited to a term proportional to $\partial_{\eta}(a^{2}(p+\epsilon))$ but that term can never been totally removed. We choose to define $\tilde{T}_{00}$ such that right-hand side is proportional to $\partial_{\eta}(a^{2}(p+\epsilon))$. It is a physical effect of non-conservation of the local stress-energy tensor due to the evolution of the background cosmological fluid.
@@
    This reproduces previous results [deVega:1998ia, Ashtekar:2015lxa, Date:2015kma, Compere:2023ktn].
@@
    $(\square+\frac{\alpha(\alpha+1)}{\eta^{2}})\chi_{0i}$ $=-16\pi Ga\delta\tilde{T}_{0i},$ (2.31b)
@@
    where $z_{\text{emiss}}=1/a(t_{\text{emiss}})-1$ is the redshift at emission time. This approximation is extremely well obeyed for all expected sources up to redshift $z\sim 100$ at the Einstein Telescope ($f_{\text{gw}}\approx 10-2000$ Hz) [ET:2025xjr] and LISA ($f_{\text{gw}}\approx 10^{-4}-1$ Hz) [LISAConsortiumWaveformWorkingGroup:2023arg] as $\varepsilon\sim 10^{-19}$ and $\varepsilon\sim 10^{-14}$, respectively. For such sources and detectors, our relativistic formalism is unnecessary, as the geometric optics approximation can be used.
@@ 3 Green’s function for power law cosmologies
where $\delta_{4}(x,x^{\prime})$ is the invariant Dirac functional [Poisson:2011nh]. Following the Hadamard construction of the Green’s function, we choose the ansatz
@@
and later we will recover the retarded Green’s function by taking the limit $\epsilon\to 0^{+}$. Using the relation $g^{\alpha\beta}\nabla_{\alpha}\sigma\nabla_{\beta}\sigma:=g^{\alpha\beta}\sigma_{\alpha}\sigma_{\beta}=2\sigma$, and the distributional identities [Poisson:2011nh]
@@
In the limit $\epsilon\to 0^{+};\epsilon\delta^{\prime}(\sigma+\epsilon)\to 0$, $\epsilon\delta^{\prime\prime}(\sigma+\epsilon)\to 2\pi\delta_{4}(x,x^{\prime})$ [Poisson:2011nh], Eq. (3.6) becomes
@@
These two equations determine $U(x,x^{\prime})$ uniquely as [Poisson:2011nh]
@@
We have to look for a solution of the characteristic initial value problem that satisfies both Eqs. (3.14) and (3.19) simultaneously. This problem can be solved by a series [Friedlander:2010eqa] [^5] We thank Abraham Harte for pointing out this reference.
@@
The Green’s function of matter dominated universe was studied by Waylen [Waylen:1978dx, a2540274-da7b-3e43-b5e5-9617401299cd] using the Hadamard construction. However, his analysis used a formulation of linearized Einstein’s equations which is valid for a maximally symmetric background but which is incorrect for an FRLW background, see his Eq. (4) of [Waylen:1978dx]. Later on, the Green’s functions were studied in [Caldwell:1993xw] for a generic FLRW universe. However, even for transverse, traceless, synchronous perturbations (which actually cannot be obtained as gauge conditions in the presence of sources), our perturbed equation in Eq. (2.8) does not match with that of Eq. (3.3) in [Caldwell:1993xw]. Our Green’s functions therefore disagree with these results.
The Green’s function of Eq. (3.1) has been obtained in [Haas:2004kw] in the context of the propagation of a scalar charge in power-law cosmologies. In [Haas:2004kw], a generic solution in the case $C=l(l+1)$ with $l$ integer for $V(x,x^{\prime})$ was obtained in terms of the Appell $F_{4}$ function of two variables,
@@
The Green’s function was constructed by Chu in Appendix B of [Chu:2016qxp] and derived for power law cosmologies in Eq. (116) of [Chu:2016ngc]. For the wave operator with $C=\alpha(\alpha\pm 1)$, we find that we can rewrite the Green’s function (3.36) as
@@
where $P_{\pm\alpha}[x]$ are the Legendre polynomials. This exactly matches with [Chu:2016qxp, Chu:2016ngc] after switching to mostly minus convention. We therefore confirm the Green’s functions of Chu [Chu:2016qxp, Chu:2016ngc] with our independent derivation.
For completeness, we note that the Green’s function was also computed in the Fourier domain in [Faraoni:1991xe, Iliopoulos:1998wq] in terms of Hankel functions. Since our derivation is in position space, we cannot directly compare our expressions with their results.
@@ 4 Multipolar decomposition of the stress-energy tensor
We aim to study compact sources which are characterized by their most relevant multipolar components. Before solving the inhomogeneous wave equations, we will start by generalizing the multipolar decomposition of the stress-energy tensor around a de Sitter background [Compere:2023ktn] to the case of a generic spatially flat FLRW background. We will then be able to perform a quadrupolar truncation.
@@ 4.1 Conservation equations
These equations have the same form as previously derived around de Sitter, see Eqs. (2.37)-(2.38) of [Compere:2023ktn] except that in FRLW universe, one has generically an additional source term proportional to $s(\eta)$. We define the angular momentum (or odd parity dipole moment) as
@@
$S_{ij}(\eta)=\int d^{3}x^{\prime}a(\eta)T_{ij}(\eta,x^{\prime})=\frac{1}{2}\partial_{t}\left(\partial_{t}Q_{ij}^{(\rho)}-2HQ_{ij}^{(\rho)}+HQ_{ij}^{(p)}-s\hat{\chi}_{ij}\right).$ (4.13)
@@
which generalizes Eq. (2.50) of [Compere:2023ktn] obtained for a de Sitter background to a generic spatially flat FRLW background.
@@ 4.2 Consistent quadrupolar truncation
In all subsequent sections, we will restrict our analysis to the quadrupolar order, which captures the dominant effects of gravitational radiation. The quadrupolar truncation was consistently defined in [Compere:2023ktn]. It amounts to fixing
@@
The equation of motion (2.28) then implies that $\chi_{L}=0$ for all $\ell>2$. All definitions provided in [Compere:2023ktn] in the case of de Sitter background then extend straightfowardly to the case of a background FRLW spacetime.
@@ 5.3 Tensor inhomogeneous solutions
Due to the fact that the Hubble “constant” is varying in matter dominated cosmologies, the tail integral is non-trivial. Gravitational waves propagate inside the light-cone in these spacetimes. This behavior differs from gravitational waves in de Sitter spacetime where the tail integral reduced to an instantenous term and a boundary term at the past cosmological horizon [Compere:2023ktn].
@@
For de Sitter the tail integral can be performed exactly and it reduces to an instantaneous term and a boundary term on the initial time slice $\eta=\eta_{i}$ as q=1q=-1 and the second term in Eq. (5.38) vanishes. The solution (5.40) exactly matches with the de Sitter solution obtained in Eq. (2.92) of [Compere:2023ktn] after neglecting the boundary terms at the past cosmological horizon. This provides a non-trivial cross-check of our solution (5.32).
@@ 6 Conclusion
Assuming power law cosmologies, we derived the Green’s functions for the two relevant potentials, which contain a light-cone part and a tail part, in terms of a hypergeometric function. We found that it matches a Green’s function derived by Chu [Chu:2016qxp, Chu:2016ngc], and compared our results with other proposals [Waylen:1978dx, a2540274-da7b-3e43-b5e5-9617401299cd, Haas:2004kw]. Using the consistent quadrupolar truncation defined in [Compere:2023ktn], we derived the metric perturbation up to quadrupolar order for an arbitrary nearly-compact source. This metric perturbation not only includes the propagating GW but also the scalar potentials that are generated by the nearly-compact source. We emphasize that our formalism applies beyond the geometric optics approximation and is therefore relevant for the study of perturbations produced by primordial sources.

### v2

@@ 1 Introduction
For all sources that obey this condition, the effect of the cosmological background on gravitational wave propagation simply amounts to a redshift of the GW frequency, the modulation of the GW amplitude by the scale factor and a redshift of the masses of the source [1, 2].
@@
The main formalism currently used for the computation of perturbations of FRLW spacetimes is based on the mathematically elegant SVT decomposition, which allows to fully decouple scalar, vector and tensor modes and write down gauge invariant cosmological perturbations [3, 4, 5]. When combined with inflation and quantum mechanics, cosmological perturbation theory allows to infer the large scale structure of the current universe [6].
Due to the spatial non-locality of the SVT decomposition, perturbations are typically expanded in Fourier modes. Fourier domain is convenient for the study of the fundamental stochastic GW background as the spectrum of perturbations generated during inflation is nearly scale invariant [7]. However, when considering individual localized sources, the SVT formalism requires to solve Poisson equations in order to reconstruct the metric perturbation, which is not straightforward. It is convenient instead to derive the perturbation equations in position space as most of the tools derived in post-Newtonian methods in the non-cosmological context are precisely derived in position space using a basis of symmetric trace-free (STF) tensors [8]. In order to fill this gap, we derive in this paper a formalism that solves for the propagation of GW in FRLW universe without using neither the geometric optics approximation nor the SVT decomposition. Instead, the equations will be decoupled in a basis of symmetric trace-free (STF) tensors by introducing a generalized harmonic gauge. Our formalism is therefore adapted to the study of compact sources in the early universe, such as e.g. cosmic strings or primordial black hole mergers [9]. The generalized harmonic gauge that we define is the first gauge to our knowledge that allows to fully decouple the linearized perturbations without using a SVT decomposition, which departs from standard treatments of cosmological perturbation theory [10].
@@
One important question in cosmology is the very definition of a compact source. We emphasize in our treatment that defining a vanishing stress-energy tensor perturbation outside a compact support is fundamentally inconsistent with the local conservation of the stress-energy tensor. This will lead us to define a nearly spatially compact source, which still admits a non-compact energy density perturbation. We will finally derive the explicit metric perturbations for a nearly compact source in power law cosmologies. We will follow the consistent quadrupolar truncation scheme introduced in [16].
@@ 2 Decoupled perturbations in generalized harmonic gauge
Let us first specialize to the de Sitter background where $\Lambda=3H^{2}$ and $a=-\frac{1}{H\eta}$, which implies $(\frac{\dot{a}}{a})=-\frac{1}{\eta}$ and q=1q=-1. In that case, the gauge condition reduces to $\bar{\nabla}_{\alpha}\tilde{h}^{\alpha}_{\;\;\mu}=\frac{2\Lambda}{3}\eta\tilde{h}_{0\mu}$, which is identical to the gauge condition defined in [22]. Around de Sitter, Eq. (2.11) becomes
@@
This equation matches exactly with Eq. (69) of [23], and Eq. (3.23) of [22] after correctly identifying the field variables. It was shown in [22] that these linearized equations can be totally decoupled after using an appropriate choice of field variables.
@@
We emphasize that unlike standard cosmological perturbation theory, we have not performed the scalar, vector, tensor (SVT) decomposition of the field variables, which is spatially non-local on constant time sections. Here, we performed a local decomposition into irreducible representations of $SO(3)$, a STF (symmetric tracefree) decomposition. The SVT decomposition in cosmological perturbation theory has found vast applications to non-localized sources in stochastic waves of astrophysical or cosmological origin. Our formalism is suited to solve for inhomogeneous solutions in terms of Green’s functions, as we will develop in Section 3. In particular, one can relate the linearized field in terms of the multipolar structure of spatially compact sources, as performed around Minkowski spacetime [8].
@@
Let us first discuss the gauge transformation of the shifted stress-energy tensor under an arbitrary linearized diffeomorphism $x^{\mu}\to x^{\mu}-\lambda\xi^{\mu}$. Using $\delta_{\xi}h_{\mu\nu}=\mathcal{L}_{\xi}\bar{g}_{\mu\nu}$ and $\delta_{\xi}\delta T_{\mu\nu}=\mathcal{L}_{\xi}\bar{T}_{\mu\nu}$, the components of shifted stress-energy tensor transform as
$\delta_{\xi}(\delta\tilde{T}_{00})$ $=$ $a^{2}(\epsilon+p)(\dot{\xi}^{0}-2aH\xi^{0}),$ (2.28)
$\delta_{\xi}(\delta\tilde{T}_{0i})$ $=$ $a^{2}(\epsilon+p)\partial_{i}\xi^{0},$ (2.29)
$\delta_{\xi}(\delta\tilde{T}_{ij})$ $=$ $\delta_{ij}\big(a^{2}\xi^{0}\dot{p}+a^{2}(\epsilon+p)(\dot{\xi}^{0}+Ha\xi^{0})\big).$ (2.30)
While the shifted stress-energy tensor is not gauge invariant, its tracefree part $\delta\tilde{T}_{\langle ij\rangle}$ is gauge invariant, and therefore a priori physical.
Importantly, one cannot absorb the right-hand side of Eq. (2.26a) by performing a local shift of the stress-energy tensor. Therefore, in the presence of a varying cosmological fluid variable $a^{2}(p+\epsilon)$, the perturbed stress-energy tensor is not conserved on its own. Using our definition of the shifted stress-energy tensor, this non-conservation can be confined to the $\eta$ component of the shifted stress-energy conservation equation, and be can limited to a term proportional to $\partial_{\eta}(a^{2}(p+\epsilon))$ but that term can never been totally removed. We choose to define $\delta\tilde{T}_{00}$ such that right-hand side is proportional to $\partial_{\eta}(a^{2}(p+\epsilon))$. It is a physical effect of non-conservation of the local stress-energy tensor due to the evolution of the background cosmological fluid.
@@
    This reproduces previous results [22, 24, 23, 16].
@@
    $(\square+\frac{\alpha(\alpha+1)}{\eta^{2}})\tilde{\chi}_{0i}$ $=-16\pi Ga\delta\tilde{T}_{0i},$ (2.34b)
@@
    where $z_{\text{emiss}}=1/a(t_{\text{emiss}})-1$ is the redshift at emission time. This approximation is extremely well obeyed for all expected sources up to redshift $z\sim 100$ at the Einstein Telescope ($f_{\text{gw}}\approx 10-2000$ Hz) [27] and LISA ($f_{\text{gw}}\approx 10^{-4}-1$ Hz) [28] as $\varepsilon\sim 10^{-19}$ and $\varepsilon\sim 10^{-14}$, respectively. For such sources and detectors, our relativistic formalism is unnecessary, as the geometric optics approximation can be used.
@@ 3 Green’s function for power law cosmologies
where $\delta_{4}(x,x^{\prime})$ is the invariant Dirac functional [29]. Following the Hadamard construction of the Green’s function, we choose the ansatz
@@
and later we will recover the retarded Green’s function by taking the limit $\epsilon\to 0^{+}$. Using the relation $g^{\alpha\beta}\nabla_{\alpha}\sigma\nabla_{\beta}\sigma:=g^{\alpha\beta}\sigma_{\alpha}\sigma_{\beta}=2\sigma$, and the distributional identities [29]
@@
In the limit $\epsilon\to 0^{+};\epsilon\delta^{\prime}(\sigma+\epsilon)\to 0$, $\epsilon\delta^{\prime\prime}(\sigma+\epsilon)\to 2\pi\delta_{4}(x,x^{\prime})$ [29], Eq. (3.6) becomes
@@
These two equations determine $U(x,x^{\prime})$ uniquely as [29]
@@
We have to look for a solution of the characteristic initial value problem that satisfies both Eqs. (3.14) and (3.19) simultaneously. This problem can be solved by a series [30] [^5] We thank Abraham Harte for pointing out this reference.
@@
The Green’s function of matter dominated universe was studied by Waylen [31, 32] using the Hadamard construction. However, his analysis used a formulation of linearized Einstein’s equations which is valid for a maximally symmetric background but which is incorrect for an FRLW background, see his Eq. (4) of [31]. Later on, the Green’s functions were studied in [20] for a generic FLRW universe. However, even for transverse, traceless, synchronous perturbations (which actually cannot be obtained as gauge conditions in the presence of sources), our perturbed equation in Eq. (2.8) does not match with that of Eq. (3.3) in [20]. Our Green’s functions therefore disagree with these results.
The Green’s function of Eq. (3.1) has been obtained in [21] in the context of the propagation of a scalar charge in power-law cosmologies. In [21], a generic solution in the case $C=l(l+1)$ with $l$ integer for $V(x,x^{\prime})$ was obtained in terms of the Appell $F_{4}$ function of two variables,
@@
The Green’s function was constructed by Chu in Appendix B of [14] and derived for power law cosmologies in Eq. (116) of [17]. For the wave operator with $C=\alpha(\alpha\pm 1)$, we find that we can rewrite the Green’s function (3.36) as
@@
where $P_{\pm\alpha}[x]$ are the Legendre polynomials. This exactly matches with [14, 17] after switching to mostly minus convention. We therefore confirm the Green’s functions of Chu [14, 17] with our independent derivation.
For completeness, we note that the Green’s function was also computed in the Fourier domain in [33, 34] in terms of Hankel functions. Since our derivation is in position space, we cannot directly compare our expressions with their results.
@@ 4 Multipolar decomposition of the stress-energy tensor
We aim to study compact sources which are characterized by their most relevant multipolar components. Before solving the inhomogeneous wave equations, we will start by generalizing the multipolar decomposition of the stress-energy tensor around a de Sitter background [16] to the case of a generic spatially flat FLRW background. We will then be able to perform a quadrupolar truncation.
@@ 4.1 Conservation equations
These equations have the same form as previously derived around de Sitter, see Eqs. (2.37)-(2.38) of [16] except that in FRLW universe, one has generically an additional source term proportional to $s(\eta)$. We define the angular momentum (or odd parity dipole moment) as
@@
$S_{ij}(\eta)=\int d^{3}x^{\prime}a(\eta)\delta\tilde{T}_{ij}(\eta,x^{\prime})=\frac{1}{2}\partial_{t}\left(\partial_{t}Q_{ij}^{(\rho)}-2HQ_{ij}^{(\rho)}+HQ_{ij}^{(p)}-s\hat{\chi}_{ij}\right).$ (4.13)
@@
which generalizes Eq. (2.50) of [16] obtained for a de Sitter background to a generic spatially flat FRLW background.
@@ 4.2 Consistent quadrupolar truncation
In all subsequent sections, we will restrict our analysis to the quadrupolar order, which captures the dominant effects of gravitational radiation. The quadrupolar truncation was consistently defined in [16]. It amounts to fixing
@@
The equation of motion (2.31) then implies that $\chi_{L}=0$ for all $\ell>2$. All definitions provided in [16] in the case of de Sitter background then extend straightfowardly to the case of a background FRLW spacetime.
@@ 5.3 Tensor inhomogeneous solutions
Due to the fact that the Hubble “constant” is varying in matter dominated cosmologies, the tail integral is non-trivial. Gravitational waves propagate inside the light-cone in these spacetimes. This behavior differs from gravitational waves in de Sitter spacetime where the tail integral reduced to an instantenous term and a boundary term at the past cosmological horizon [16].
@@
For de Sitter the tail integral can be performed exactly and it reduces to an instantaneous term and a boundary term on the initial time slice $\eta=\eta_{i}$ as q=1q=-1 and the second term in Eq. (5.38) vanishes. The solution (5.40) exactly matches with the de Sitter solution obtained in Eq. (2.92) of [16] after neglecting the boundary terms at the past cosmological horizon. This provides a non-trivial cross-check of our solution (5.32).
@@ 6 Conclusion
Assuming power law cosmologies, we derived the Green’s functions for the two relevant potentials, which contain a light-cone part and a tail part, in terms of a hypergeometric function. We found that it matches a Green’s function derived by Chu [14, 17], and compared our results with other proposals [31, 32, 21]. Using the consistent quadrupolar truncation defined in [16], we derived the metric perturbation up to quadrupolar order for an arbitrary nearly-compact source. This metric perturbation not only includes the propagating GW but also the scalar potentials that are generated by the nearly-compact source. We emphasize that our formalism applies beyond the geometric optics approximation and is therefore relevant for the study of perturbations produced by primordial sources.